#36    A research report states that there is a significant difference between treatments for an independent-measures design with 28 degrees of freedom, t = 2.27.

How many individuals participated in this study?
Answer #1

here the degrees of freedom = 28

degrees of freedom = n -1

n = df + 1

n = 28 + 1 = 29

Hence 29 individuals participated in this study